solve by the elimination method:
5x+6y=4
-10x-12y=-8
1
Expert's answer
2011-04-12T09:19:53-0400
& 5x + 6y = 4 |x(-2)
-10x - 12y = -8;

-10x - 12y = -8
-10x - 12y = -8

There are infinite number of solutions, x and y& are related as y = -5/6 x - 2/3.